OpenStudy (anonymous):

How To Factor (a-2b)^2 - (a-5)^2

OpenStudy (raden):

hint : the factor of m^2 - n^2 = (m + n)(m - n) now given (a-2b)^2 - (a-5)^2 it mean m = a-2b and n = a-5 applies formula above to make clear